• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Punjabi

    Log entry time 07:58:06 on January27,2002

    Entry number 80985

    keyword=Shift Summary

    Started shift in kinematics 4G.
    00:00 Changed target to 15 cm LH2 with radiator. Had to reboot target IOC twice.
    Changed left spectrometer to 18.21 deg. nd momentum to 3.742.

    12:33 Sarted run 1747. First Production run for 4G kinematics.

    1:42 finished run 1747,1748,1749 and 1750.

    1:42 CODA problems. ROC4 was hanging. Could not reset it. Kcoda 3 times. Finally coda started.

    2:05 Started run 1754. Run 1751,52 and 53 do not exist.

    3:12 Finished run 1754, 1755, 1756 and 1757.

    3:13 Started run 1758.

    4:21 Finished run 1758,1759,1760,1761.

    4:22 Started run 1762.

    5:06 Finished run 1762,1763.

    5:06 Stopped run 1764. MCC called. Beam off. Vacuum problem. They called experts. Estimate 30 minuites to an hour.

    7:00 beam is still off. Tech is in. May be another 30 to 45 minuites.

    8:00 Beam still off.

    8:00 Dipole and Q1 went down. Mark Stevens will be here. Will take controlled access to fix it.

    8:00 We collected 0.26 C at 4G kinematics.